You must state your answer in words. You must set up a system of two equations in two variables.
You must state what the variables represent.
You must show all the steps in solving the system of equations.
An express train and a local train both leave Gray’s Lake at 12:00 noon and head for Chicago 60 miles away. The express travels twice as fast as the local and arrives 2 hours ahead of it. Find the speed of each train.
Thank You!

Answer by stanbon(57224)

An express train and a local train both leave Gray’s Lake at 12:00 noon and head for Chicago 60 miles away. The express travels twice as fast as the local and arrives 2 hours ahead of it. Find the speed of each train.
--------------------------
Express DATA:
distance = 60 miles ; rate = E mph ; time = 60/E hrs
-------------------
Local DATA:
distance = 60 miles ; rate = L mph ; time = 60/L hrs
--------------------
Equations:
E = 2L
60/L - 60/E = 2
------------------
Substitute to solve for L:
60/L - 60/2L = 2
Multiply thru by 2L to get:
120 - 60 = 4L
60 = 4L
L = 15 mph (speed of the local train)
E = 2L = 30 mph (speed of the express train)
================================================
Cheers,
Stan H.